OpenStudy (anonymous):

Square root of x-5 = 2square root of 6 .

OpenStudy (anonymous):

\[\sqrt{x-5}=2\sqrt{6}\] \[x-5=24\] \[x=29\]

OpenStudy (anonymous):

step one, square both sides step to, add 5 to both sides